given that (-3,-5) and (-1,-4) are points on the same line find a point-slope form of the equation for this line i know that I have to find the slope first which is y1-y2/x1-x2 and that formula is -5--4/-3--1 so the answer to that part of the equation is 1/2 from there I am lost as to how to write it in point-slope form
Answer by rothauserc(4718)

both points are on the same line so you can pick either one
:
point slope form is
:
(y-y1) = m(x-x1) where m is the slope
:
let's pick the first point and substitute for x1 and y1
:
******************
(y+5) = (x+3)/2
******************
: